OpenStudy (anonymous):

3x + y = 11 –3x + 2y = 1 2 4 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (lopezking1):

4

OpenStudy (anonymous):

Im not quite sure what the sentences are about... but if you add the equations you get 3y = 12 which means y =4, then you back plug and get 3x+4= 11 so 3x = 7, x = 7/3... and then you can check and -3(7/3) + 2y = 1 and it works out. so y=4, x=7/3..

OpenStudy (anonymous):

thanks